OpenStudy (anonymous):

Imagine any number from 1 to 9, Multiply it with \(\color{blue}{9}\), Then, Multiply it with \(\color{green}{12345679}\), Now, check what you got.. \(\checkmark\)..

OpenStudy (anonymous):

Just a normal LKG Trick.. :)

OpenStudy (unklerhaukus):

\[\color{teal}8\times\color{blue}9\times\color{green}{12345679}=\color{teal}{888888888}\]

OpenStudy (anonymous):

It is showing your Imagined number the times you multiply it with..

OpenStudy (unklerhaukus):

can you explain why this works?

OpenStudy (anonymous):

I have posted it for you, so that you explain us why this works.. :P Now explain.. :)

OpenStudy (anonymous):

You were writing, no??

OpenStudy (anonymous):

Don't tell me you are still writing uncle rocks??

OpenStudy (unklerhaukus):

oh but i am

OpenStudy (anonymous):

Are you sure you are writing at this post only?? :P

OpenStudy (anonymous):

@UnkleRhaukus don't use latex otherwise, it would take 7-8 hours.. :)

OpenStudy (unklerhaukus):

\[\color{blue}9\times\color{green}{12345679}\\=90000000+18000000+2700000+360000+45000+5400+630+81\\=\qquad108000000\qquad+\qquad3060000\quad+\quad\qquad50400\qquad+711\\=\qquad\qquad\qquad\qquad 111060000\qquad\qquad+\qquad\qquad\qquad51111\\=\qquad\qquad\qquad\qquad\qquad\qquad\qquad111111111\]

OpenStudy (anonymous):

He he he.. This one I know that 9*12345679 will make to 111111111

OpenStudy (unklerhaukus):

still dosent make much sense to me

OpenStudy (anonymous):

For your effort, that you wrote that much, A Honor Of Writing I am awarding to you..

OpenStudy (anonymous):

It works with 0 too, if you imagine the right number of zeroes \[0\times9\times12345679=000000000\] :)